OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (mayankdevnani):

2x+12-10=12+4x-4 2x+2=8+4x 2x-4x=8-2 -2x=6 -x=3 x=-3

OpenStudy (mayankdevnani):

ok @bethedw20

OpenStudy (anonymous):

thank you